OpenStudy (anonymous):

Find the ordered pair that represents the vector from B(–6, 2) to C(–12, 7). A. BC= (1, -10) B. BC=(-18, 9) C. BC= (2, -5) D. BC= (-6, 5) Find the magnitude of vector WX for W(–3, 5, 4) and X(9, 5, 3). A. Square Root of 137 B. Square Root of 145 C.140 D. 137

OpenStudy (raden):

vector BC = c - b = (-12-(-6), 7-2) = (... , ...)

OpenStudy (anonymous):

So what do you do? Take -12 - 7 or take both coordinates of b and minus them from all coordinates of c?

OpenStudy (raden):

-12 - (-6) = -12+6 = ... ? im not sure u cant calculate of 7-2 = ... ? then represent as a coordinat, or it can represented as a matrice

OpenStudy (anonymous):

I don't know how to do this at all, I'm sorry. Uh, ? I'm supposed to find the ordered pair that represents the vector from B & C. . . you're confusing me even more with the matrices and stuff.

OpenStudy (anonymous):

vector WX = <12,0,-1> its magnitude is sqwr(12^2 +(-1)^2 = sqr (145)

OpenStudy (anonymous):

Alright, now that makes sense, thank you.

OpenStudy (anonymous):

to find out a vector from 2 points, you take <x2-x1, y2 -y1>, pay attention to the notation < >, it show that is a vector, not a point

OpenStudy (anonymous):

and then, from that information, you just take square root of the x ^2 + y^2 (x, y from the vector you 've just found out. by that way, you can find out any vector, and its magnitude no matter what it is in R^2 or R^3 Hope this help

OpenStudy (anonymous):

Thank you very much, the both of you. I'm apologize, when it comes to math - it takes a lot to get me to understand what I'm supposed to be doing.